Two trains are on the same track going in opposite directions. One is traveling at 55 mph and the other is traveling at 45 mph and they are 250 miles apart. How long until they hit?

Answer by tommyt3rd(5050)

Speed:
s=d/t
Trains moving toward each other means their speeds add - so we write:
t=d/s=250/(55+45)=2.5

2 and 1/2 hours

:)